Suppose you currently live and work in​ Cleveland, earning a salary of ​$70 comma 000 per year and spending ​$20 comma 000 for housing. You just heard that you will be transferred to a city in California where housing is 50 percent more expensive. In negotiating a new​ salary, your objective is to keep your real income constant.  
Your new target salary is ​$
nothing. ​(Enter your response rounded to the nearest dollar.​)
1
Expert's answer
2020-04-06T10:13:40-0400

New housing cost"=20000\u00d71.5=" $"30000;"

permanent income "=70000-20000=" $"50000;"

New salary"=50000+30000="$ "80000;"